# Andreas Schlaginhaufen SYCAMORE, EPFL &Reda Ouhamma SYCAMORE, EPFL &Maryam Kamgarpour

### v1

@@ 1 Introduction
Reinforcement learning (RL) is a fundamental paradigm in machine learning, where agents learn to make sequential decisions by interacting with an environment to maximize cumulative rewards (Barto, 2021). RL has enabled advances in domains such as game play (Silver et al., 2017), robotics (Todorov et al., 2012), or autonomous driving (Lu et al., 2023). However, the practicality of RL is hindered by the challenge of designing rewards: crafting a reward function that aligns with human objectives is often difficult, and a misspecified reward function can lead to suboptimal or unsafe behavior (Amodei et al., 2016; Hadfield-Menell et al., 2017). This motivates the development of principled alternatives to manual reward design.
@@
Rather than relying on manually specified reward functions, reinforcement learning from human feedback (RLHF) guides learning through preference feedback: at each step, a human oracle compares trajectories and indicates which is preferable (Christiano et al., 2017). This preference signal is often much easier to provide than engineering a reward function (Pereira et al., 2019; Lee et al., 2023). RLHF has proven to be effective in robotics (Jain et al., 2013) and, more recently, fine-tuning of large language models (Stiennon et al., 2020; Ziegler et al., 2019; Rafailov et al., 2023). This highlights the practical relevance of RLHF compared to reward-based learning.
@@
Despite its empirical success, the theoretical foundations of RLHF are still in development. Existing works first studied the simpler setting of dueling bandits. In this context, the learner selects pairs of actions and observes noisy preference feedback (Yue et al., 2012; Komiyama et al., 2015). Classical algorithms for regret minimization in this setting include approaches based on zeroth-order optimization (Yue and Joachims, 2009) or the principle of optimism (Ailon et al., 2014). A key challenge in this setting is reducing the number of preference queries. For this purpose, several recent works propose strategic query selection strategies for dueling bandits (Das et al., 2024; Liu et al., 2024; Scheid et al., 2024; Mukherjee et al., 2024), often hinging on optimal experimental design mechanisms (Pukelsheim, 2006). However, such approaches are usually limited to finite-armed bandits, where the resulting optimization problems can be solved efficiently.
@@
In the online RL setting, the theory of RLHF has received increasing attention, with several works establishing either regret or probably approximately correct (PAC) guarantees. PAC-RL methods aim to identify a near-optimal policy with high probability (Xu et al., 2020; Novoseller et al., 2020; Zhu et al., 2023), while regret-based approaches provide bounds on the cumulative reward during learning (Pacchiano et al., 2021; Chen et al., 2022; Wu and Sun, 2023). Similar to dueling bandits, a central challenge in RLHF is to actively select informative trajectory comparisons to drive learning. In RL, however, this active-learning problem presents additional difficulties: First, the learner cannot freely choose arbitrary state-action pairs or trajectories, but must reach them through exploration (Wagenmaker et al., 2022). Second, many existing approaches with guarantees (Pacchiano et al., 2021; Zhan et al., 2024) rely on maximizing an exploration bonus involving a norm of state distributions – a problem which is known to be computationally intractable even in tabular settings (Efroni et al., 2021).
@@
To sidestep these challenges, another line of work focuses on RLHF with offline data. In this setting, learning proceeds over a fixed pre-collected dataset of trajectory preferences (Zhu et al., 2023; Zhan et al., 2023). Although this offline paradigm avoids the need for online exploration and active query selection, it depends critically on having access to sufficiently diverse and informative preference data a priori (Rashidinejad et al., 2021; Xie et al., 2021; Zanette et al., 2021; Zanette, 2023) — a requirement that can be difficult to meet in practice. Hence, this merely shifts the exploration and active learning challenges to the data collection phase.
@@
Despite progress on statistical guarantees in RLHF, two key challenges remain open: the design of tractable algorithms for active preference query selection and reducing the workload of human preference annotators. In existing approaches, a human must provide feedback at every round, which is impractical in real-world applications. Our goal in this work is to develop RLHF algorithms that are computationally efficient, reduce the demand for human feedback, and actively select informative queries. Some recent work has made progress on tractability. For instance, Wu and Sun (2023) proposes a randomized exploration algorithm with regret guarantees, but their method is limited to linear dynamics. In parallel, Wang et al. (2023) introduces a general reduction from RLHF to standard RL and establishes PAC-style guarantees under RL oracle access. However, neither approach addresses the open challenges of reducing feedback requirements or enabling active query selection.
@@
In this work, we focus on reinforcement learning from human feedback (RLHF) and develop meta-algorithms that reduce the RLHF problem to standard RL by leveraging existing RL algorithms as subroutines. Our approach combines randomized exploration for tractability, lazy updates to reduce human workload, and experimental design to actively select informative preference queries. We provide both regret and PAC-style guarantees under RL oracle assumptions. Our contributions are as follows:
  • We propose general meta-algorithms for RLHF using RL oracles, and provide provable regret and PAC guarantees in general MDPs.
  • We present a second meta-algorithm with better scalability and query efficiency thanks to: Lazy updates, inspired by linear bandits (Abbasi-Yadkori et al., 2011), which enables parallelization of the preference oracle calls; Greedy optimal design, which selects high-quality preference queries and improves sample efficiency.
  • We provide empirical results showing that: Our algorithm is implementable and competitive with reward-based RL; The improved algorithm achieves comparable performance while significantly reducing the query complexity. [^1] The code is openly accessible at https://github.com/andrschl/isaac_rlhf.
@@ 2 Preliminaries
[…]
@@
We consider an infinite-horizon [^2] Our results extend directly to the finite-horizon setting as well. However, we focus on the infinite-horizon discounted setting, as it is more commonly encountered in deep reinforcement learning. Markov decision process (MDP) =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\nu_{0},P,r,\gamma\}$ with compact state and action spaces $\mathcal{S}\subset\mathbb{R}^{n}$ and $\mathcal{A}\subset\mathbb{R}^{m}$, respectively, initial state distribution $s_{0}\sim\nu_{0}$, transition law $s_{h+1}\sim P(\cdot|s_{h},a_{h})$, and discount rate $\gamma\in(0,1)$. We assume a linear reward model $r_{\theta^{*}}(s,a):=\langle\theta^{*},\phi(s,a)\rangle$, where $\left\lVert\theta^{*}\right\rVert\leq B$ and $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature mapping such that $\max_{s,a}\left\lVert\phi(s,a)\right\rVert\leq L$. We denote the set of all trajectories as $\mathcal{T}:={(\mathcal{S}\times\mathcal{A})}^{\infty}$, and the distribution over $\mathcal{T}$ induced by a stationary Markov policy $\pi:\mathcal{S}\to\Delta_{\mathcal{A}}$ as $\mathbb{P}_{\pi}$. For a trajectory $\tau=(s_{0},a_{0},s_{1}\ldots)\in\mathcal{T}$, we denote the discounted sum of features by $\phi(\tau):=\sum_{h=0}^{\infty}\gamma^{h}\phi(s_{h},a_{h})$ and the feature expectation of a policy $\pi$ by $\phi(\pi):=\mathbb{E}_{\tau\sim\mathbb{P}_{\pi}}\left[{\phi(\tau)}\right]$. Furthermore, given a reward parameter $\theta$, we denote the value of a policy $\pi$ by $V_{\theta}^{\pi}:=\mathbb{E}_{\tau\sim\mathbb{P}_{\pi}}\left[{\sum_{h=0}^{\infty}\gamma^{h}r_{\theta}(s_{h},a_{h})}\right]=\langle\theta,\phi(\pi)\rangle$ and the optimal value by $V_{\theta}^{*}:=\max_{\pi}V_{\theta}^{\pi}$.
@@
This preference model is a special case of the Plackett-Luce model (Plackett, 1975; Luce et al., 1959), and is commonly used in the dueling bandit setting as well as the RLHF framework (Yue et al., 2012; Christiano et al., 2017; Ouyang et al., 2022). In particular, this model satisfies the strong stochastic transitivity property (Strzalecki, 2025). That is, for any three trajectories $\tau_{1}\succ\tau_{2}\succ\tau_{3}$ we have $\mathcal{P}(\tau_{1}\succ\tau_{3})>\max\{\mathcal{P}(\tau_{1}\succ\tau_{2}),\mathcal{P}(\tau_{2}\succ\tau_{3})\}$, capturing the monotonicity of preferences.
@@
In practice, we cannot compare trajectories of infinite length. Fortunately, many environments terminate in finite time, and otherwise one may truncate each trajectory at horizon $H=\mathcal{O}(\log_{\gamma}(\varepsilon))$, introducing at most an $\varepsilon$ error in value estimates (see e.g. (Schlaginhaufen and Kamgarpour, 2024)). For simplicity, however, we omit this truncation step in our presentation.
@@
Cumulative regret has been widely adopted in the RL and RLHF literature (Abbasi-Yadkori et al., 2011; Zanette et al., 2020; Wang et al., 2023; Zhan et al., 2024). However, cumulative regret doesn’t provide us with a guarantee of the last iterate’s suboptimality. As a second metric, we therefore also consider the suboptimality of an output policy.
@@
Suboptimality has previously been considered as a performance metric for offline RLHF Zhu et al. (2023), contextual bandits Das et al. (2024), and online RLHF (Wang et al., 2023). In the following, we propose a meta-algorithm that features two variants: one with theoretical guarantees on cumulative regret, and another specifically ensuring a bound on last-iterate suboptimality.
@@ 3.1 Algorithm
Our algorithm proceeds in three essential steps. First, the preference feedback data is used to estimate the reward parameter using maximum likelihood estimation. Then, a reward parameter is sampled from a Gaussian distribution, which is reminiscent of linear Thompson sampling (Abeille and Lazaric, 2017). Finally, an RL oracle is used to find an approximately optimal policy for the sampled reward parameter, and we query the preference oracle by comparing one trajectory from this new policy against one from the previous policy.
@@
Considering our preference model (1), a standard approach for estimating the reward parameter $\theta^{*}$ is via maximum likelihood estimation. Given a pair of trajectories $\tau_{k}=(s_{h,k},a_{h,k})_{h=0}^{\infty}$ and $\tau_{k}=(s^{\prime}_{h,k},a^{\prime}_{h,k})_{h=0}^{\infty}$ we consider the design points $x_{k}:=\phi(\tau_{k})-\phi(\tau^{\prime}_{k})=\sum_{h=0}^{\infty}\gamma^{h}(\phi(s_{h,k},a_{h,k})-\phi(s^{\prime}_{h,k},a^{\prime}_{h,k}))$ and the preference labels $y_{k}=\mathbbm{1}(\tau_{k}\succ\tau^{\prime}_{k})$. In round $t$, the preference dataset is $\mathcal{D}_{t}=\{(x_{k},y_{k})\}_{k=1}^{t-1}$ and the corresponding (constrained) maximum likelihood estimator (MLE) is given by $\hat{\theta}_{t}=\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t}}(\theta)$, where
@@
is the negative log-likelihood of the Bradley-Terry model (1). The loss function (2) is the familiar logistic loss from logistic regression (Shalev-Shwartz and Ben-David, 2014). In particular, it is a convex problem that can be solved efficiently using standard methods such as LBFGS (Liu and Nocedal, 1989). Moreover, we have the following time-uniform confidence result.
Lemma 3.1.
Let $\lambda>0$ and define the design matrix at time $t$ given by $V_{t}=\lambda I+\sum_{s=1}^{t-1}x_{t}x_{t}^{\top}$. Then, with probability $1-\delta$, for all $t\geq 1$, the true reward parameter $\theta^{*}$ is contained in the ellipsoid
$\mathcal{E}_{t}(\delta):=\left\{{\theta:\left\lVert\theta-\hat{\theta}_{t}\right\rVert_{V_{t}}^{2}\leq\beta_{t}(\delta)^{2}:=\mathcal{O}\left({\kappa\left[{\log\left({\frac{1}{\delta}}\right)+d\log\left({\frac{t}{d}}\right)}\right]+\lambda}\right)}\right\}.$
Here, $\kappa:=\max_{\theta\in\mathcal{B}^{d}(B),x\in\mathcal{B}^{d}(2LH_{\gamma})}1/\dot{\sigma}\left({\langle\theta,x\rangle}\right)$ denotes the Lipschitz constant of the inverse sigmoid function, and $H_{\gamma}=(1-\gamma)^{-1}$ the effective horizon of the MDP.
The above lemma hinges on a likelihood-ratio confidence set from Lee et al. (2024). The proof and the precise constants are deferred to Appendix A.
@@
Compared to the standard analysis of stochastic linear bandits Abbasi-Yadkori et al. (2011), our parameter $\beta_{t}$ includes an additional factor of $\sqrt{\kappa}$, which arises naturally due to preference-based feedback. This result improves upon the bound provided by Zhu et al. (2023), which incurs a larger factor of $\kappa$ instead of $\sqrt{\kappa}$. While the $\sqrt{\kappa}$ factor can theoretically be avoided by constructing confidence sets using the Hessian of the negative log-likelihood $\mathcal{L}_{\mathcal{D}_{t}}$ (Lee et al., 2024; Das et al., 2024), it reappears in the regret bounds as shown in (Das et al., 2024). Thus, we adopt confidence sets based on $V_{t}$, as both $V_{t}$ and its inverse can be efficiently updated via rank-one operations, unlike Hessian-based approaches.
@@
Many approaches to regret minimization and pure exploration in RLHF often rely on maximizing an exploration bonus $\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}$. Although such methods yield provable guarantees for regret (Pacchiano et al., 2021) or last-iterate suboptimality (Das et al., 2024), they are computationally intractable in RL settings (see Appendix E.2 for a discussion). To address this, we adopt a randomized exploration scheme inspired by Thompson sampling algorithms for linear bandits. In line with Abeille and Lazaric (2017), we sample the reward parameter from an inflated version of the confidence set defined in Lemma 3.1, which produces a computationally efficient alternative to optimism-based approaches. Furthermore, we also show that this randomized strategy extends to the pure exploration setting.
@@
We assume access to an $(\varepsilon,\delta)$-PAC oracle, $\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}$, for the RL problem. That is, a polynomial-time algorithm that produces for every $\varepsilon>0$, $\delta>0$, and $\theta\in\mathbb{R}^{d}$ a policy $\pi=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\theta,\varepsilon,\delta)$ such that with probability at least $1-\delta$ we have $V^{*}_{\theta}-V_{\theta}^{\pi}\leq\varepsilon$.
This assumption of a PAC-RL oracle is satisfied in several settings, including tabular and linear MDPs (Dann et al., 2019; Ménard et al., 2021; Al Marjani et al., 2021; Wagenmaker et al., 2022; He et al., 2021). Moreover, it is relatively mild compared to stronger oracles considered in the RLHF literature (Zhan et al., 2024), such as reward-free algorithms (Wang et al., 2020; Kaufmann et al., 2021; Ménard et al., 2021). In practice, common choices for $\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}$ are policy optimization methods such as proximal policy optimization (PPO) (Schulman et al., 2017) or soft actor critic (Haarnoja et al., 2018), which have shown strong empirical performance in continuous control and large-scale applications such as training large language models.
@@
Our algorithm RPO presented below comes in two variants: a variant that balances exploration and exploitation for regret minimization ($\alpha=1$), and a variant for pure exploration ($\alpha=0$). In line 4, we sample a reward parameter from a confidence set inflated by $\sqrt{d}$. Then, in line 5, we compute the policy $\pi_{t}$ using an RL oracle. Finally, in line 10, we use the new preference data to update the reward parameter using maximum likelihood.
Input: Number iterations $T$, confidence $\delta>0$, regularization $\lambda>0$, algorithm type $\alpha\in\left\{{0,1}\right\}$.
1 Set $\varepsilon=1/\sqrt{T}$ and $\delta^{\prime}=\delta/5$.
2 Initialize design matrix $V_{1}=\lambda I$, preference data set $\mathcal{D}_{1}=\emptyset$, parameter $\hat{\theta}_{1}$, and policy $\pi_{0}$.
3 for $t=1,2,\dots,T$ do
       $\tilde{\theta}_{t}\sim\mathcal{N}\big{(}\alpha\hat{\theta}_{t},\beta_{t}(\delta^{\prime})^{2}V_{t}^{-1}\big{)}$;

### v2

@@ 1 Introduction
Reinforcement learning (RL) is a fundamental paradigm in machine learning, where agents learn to make sequential decisions by interacting with an environment to maximize cumulative rewards (barto2021reinforcement). RL has enabled advances in domains such as game play (silver2017mastering), robotics (todorov2012mujoco), or autonomous driving (lu2023imitation). However, the practicality of RL is hindered by the challenge of designing rewards: crafting a reward function that aligns with human objectives is often difficult, and a misspecified reward function can lead to suboptimal or unsafe behavior (amodei2016concrete; hadfield2017inverse). This motivates the development of principled alternatives to manual reward design.
@@
Rather than relying on manually specified reward functions, reinforcement learning from human feedback (RLHF) guides learning through preference feedback: at each step, a human oracle compares trajectories and indicates which is preferable (christiano2017deep). This preference signal is often much easier to provide than engineering a reward function (pereira2019online; lee2023aligning). RLHF has proven to be effective in robotics (jain2013learning) and, more recently, finetuning of large language models (ziegler2019fine; stiennon2020learning; rafailov2023direct). This highlights the practical relevance of RLHF compared to reward-based learning.
@@
Despite its empirical success, the theoretical foundations of RLHF are still in development. Existing works first studied the simpler setting of dueling bandits. In this context, the learner selects pairs of actions and observes noisy preference feedback (yue2012k; komiyama2015regret). Classical algorithms for regret minimization in this setting include approaches based on zeroth-order optimization (yue2009interactively) or the principle of optimism (ailon2014reducing). A key challenge in this setting is reducing the number of preference queries. For this purpose, several recent works propose query selection strategies for dueling bandits (das2024active; liu2024dual; scheid2024optimal; mukherjee2024optimal), often hinging on optimal experimental design mechanisms (pukelsheim2006optimal). However, such approaches are usually limited to finite-armed bandits, where the resulting optimization problems can be solved efficiently.
@@
In the online RL setting, the theory of RLHF has received increasing attention, with several works establishing either regret or probably approximately correct (PAC) guarantees. PAC-RL methods aim to identify a near-optimal policy with high probability (xu2020preference; novoseller2020dueling; zhu2023principled), while regret-based approaches provide bounds on the cumulative reward during learning (pacchiano2021dueling; chen2022human; wu2023making). Similar to dueling bandits, a central challenge in RLHF is to actively select informative trajectory comparisons to drive learning. In RL, however, this active-learning problem presents additional difficulties: First, the learner cannot freely choose arbitrary state-action pairs or trajectories, but must reach them through exploration (wagenmaker2022beyond). Second, many existing approaches with guarantees (pacchiano2021dueling; zhan2024provablereward) rely on maximizing an exploration bonus involving a norm of state distributions – a problem which is computationally intractable even in tabular settings (efroni2021reinforcement).
@@
To sidestep these challenges, another line of work focuses on RLHF with offline data. In this setting, learning proceeds over a fixed pre-collected dataset of trajectory preferences (zhu2023principled; zhan2023provable). Although this offline paradigm avoids the need for online exploration and active query selection, it depends critically on having access to sufficiently diverse and informative preference data a priori (rashidinejad2021bridging; xie2021policy; zanette2021provable; zanette2023realizability) — a requirement that can be difficult to meet in practice. Hence, this merely shifts the exploration and active learning challenges to the data collection phase.
@@
Despite progress on statistical guarantees in RLHF, a central challenge remains open: designing tractable algorithms for active preference query selection that reduce the workload of human annotators. In existing approaches, a human must provide feedback at every round, which is impractical in real-world applications. Our goal in this work is to develop RLHF algorithms that are computationally efficient, reduce the demand for human feedback, and actively select informative queries. Some recent work has made progress on computational tractability. For instance, wu2023making propose a randomized exploration algorithm with regret guarantees limited to linear dynamics, and dwaracherla2024efficient show empirically that randomized exploration is efficient for finetuning of language models. In parallel, wang2023rlhf introduce a general reduction from RLHF to standard RL and establishes PAC-style guarantees under RL oracle access. However, neither approach addresses the open challenges of reducing feedback requirements or enabling active query selection.
@@
In this work, we focus on reinforcement learning from human feedback (RLHF) and develop meta-algorithms that reduce the RLHF problem to standard RL by leveraging existing RL algorithms as subroutines. Leveraging randomized exploration for tractable and efficient preference query selection, we provide both online algorithms with regret guarantees and a preference-free algorithm with PAC-style guarantees under RL-oracle assumptions. Our contributions are as follows:
  • We propose two meta-algorithms for RLHF using RL oracles: Algorithm 1, optimized for regret minimization, and Algorithm 2, which performs preference-free exploration and defers preference collection to a single batch at the end. For these methods, we establish regret and PAC-style guarantees, respectively, holding in general MDPs.
  • We present a second meta-algorithm for regret minimization (Algorithm 3) with better scalability and query efficiency thanks to: Lazy updates, inspired by linear bandits (abbasi2011improved), which enables parallelization of the preference oracle calls; Greedy optimal design, which selects informative preference queries and improves sample efficiency.
  • We provide empirical results showing that: Our algorithms are implementable, competitive with reward-based RL, and substantially outperform a baseline that relies solely on entropy-based exploration; Algorithm 3 achieves comparable performance to Algorithm 1 while significantly reducing the query complexity. [^1] The code is openly accessible at https://github.com/andrschl/isaac_rlhf.
@@ 2 Preliminaries
Let $\mathbb{N}$ and $\mathbb{R}$ denote the sets of natural and real numbers, respectively. We write $\left\lVert\cdot\right\rVert$ for the Euclidean norm and $\langle\cdot,\cdot\rangle$ for the standard inner product in $\mathbb{R}^{d}$. Moreover, for a positive definite matrix $A\in\mathbb{R}^{d\times d}$, we denote $\left\lVert x\right\rVert_{A}:=\sqrt{\langle x,Ax\rangle}$ for the Mahalanobis norm. Furthermore, we denote the closed Euclidean ball of radius $a>0$ by $\mathcal{B}^{d}(a)\subset\mathbb{R}^{d}$, and for a compact subset $\mathcal{X}\subset\mathbb{R}^{n}$, we denote the set of all probability measures supported on $\mathcal{X}$ by $\Delta_{\mathcal{X}}$. Finally, we use the standard notation $\mathcal{O}(n)$ and $\Omega(n)$ for asymptotic upper and lower bounds, as well as $\tilde{\mathcal{O}}(n)=\mathcal{O}(n\operatorname{polylog(n)})$ for suppressing polylogarithmic terms.
@@
We consider an infinite-horizon [^2] Our results extend directly to the finite-horizon setting as well. Markov decision process (MDP) $\mathcal{M}=\{\mathcal{S},\mathcal{A},\nu_{0},P,r,\gamma\}$ with state and action spaces $\mathcal{S}\subseteq\mathbb{R}^{n}$ and $\mathcal{A}\subset\mathbb{R}^{m}$, respectively, initial state distribution $s_{0}\sim\nu_{0}$, transition law $s_{h+1}\sim P(\cdot|s_{h},a_{h})$, and discount rate $\gamma\in(0,1)$. We assume a linear reward model $r_{\theta^{*}}(s,a):=\langle\theta^{*},\phi(s,a)\rangle$, where $\left\lVert\theta^{*}\right\rVert\leq B$ and $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ is a continuous feature mapping such that $\max_{s,a}\left\lVert\phi(s,a)\right\rVert\leq L$. We denote the set of all trajectories as $\mathcal{T}:={(\mathcal{S}\times\mathcal{A})}^{\infty}$, and the distribution over $\mathcal{T}$ induced by a stationary Markov policy $\pi:\mathcal{S}\to\Delta_{\mathcal{A}}$ as $\mathbb{P}_{\pi}$. For a trajectory $\tau=(s_{0},a_{0},s_{1},\ldots)\in\mathcal{T}$, we denote the discounted sum of features by $\phi(\tau):=\sum_{h=0}^{\infty}\gamma^{h}\phi(s_{h},a_{h})$ and the feature expectation of a policy $\pi$ by $\phi(\pi):=\mathbb{E}_{\tau\sim\mathbb{P}_{\pi}}\left[{\phi(\tau)}\right]$. Furthermore, given a reward parameter $\theta$, we denote the value of a policy $\pi$ by $V_{\theta}^{\pi}:=\mathbb{E}_{\tau\sim\mathbb{P}_{\pi}}\left[{\sum_{h=0}^{\infty}\gamma^{h}r_{\theta}(s_{h},a_{h})}\right]=\langle\theta,\phi(\pi)\rangle$ and the optimal value by $V_{\theta}^{*}:=\max_{\pi}V_{\theta}^{\pi}$.
@@
This preference model is a special case of the Plackett-Luce model (plackett1975analysis; luce1959individual), and is commonly used in the dueling bandit setting as well as the RLHF framework (yue2012k; christiano2017deep; ouyang2022training). It captures that the probability of preferring $\tau$ over $\tau^{\prime}$ is increasing in the difference between their values $\langle\theta^{*},\phi(\tau)-\phi(\tau^{\prime})\rangle$.
@@
In practice, we cannot compare trajectories of infinite length. Fortunately, many environments terminate in finite time, and otherwise one may truncate each trajectory at horizon $H=\mathcal{O}(\log_{\gamma}(\varepsilon))$, introducing at most an $\varepsilon$ error in value estimates (see e.g. , (schlaginhaufen2024towards)). For simplicity, however, we omit this truncation step in our presentation.
@@
Cumulative regret has been widely adopted in the RL and RLHF literature (abbasi2011improved; zanette2020frequentist; wang2023rlhf; zhan2024provablereward). However, cumulative regret doesn’t provide us with a guarantee of the last iterate’s suboptimality. As a second metric, we therefore also consider the suboptimality of an output policy.
@@
Suboptimality has previously been considered as a performance metric for offline RLHF (zhu2023principled), contextual bandits (das2024active), and online RLHF (wang2023rlhf). In the following, we propose a meta-algorithm that features two variants: one with theoretical guarantees on cumulative regret, and another specifically ensuring a bound on last-iterate suboptimality.
@@ 3.1 Algorithm
Both algorithms share the same key ingredients: (i) estimating the reward parameter from preference feedback using maximum likelihood estimation, (ii) sampling reward parameters from a Gaussian distribution, reminiscent of linear Thompson sampling (pmlr-v54-abeille17a), and (iii) leveraging an RL oracle to find an approximately optimal policy for the sampled reward parameter. In Algorithm 1, we apply these steps iteratively and, at each round, query the preference oracle by comparing a trajectory from the new policy to one from the previous policy to minimize regret. In contrast, Algorithm 2 performs preference-free exploration and defers all preference queries to a single batch at the end.
@@
Considering our preference model (1), a standard approach for estimating the reward parameter $\theta^{*}$ is via maximum likelihood estimation. Given a pair of trajectories $\tau_{k}=(s_{h,k},a_{h,k})_{h=0}^{\infty}$ and $\tau_{k}^{\prime}=(s^{\prime}_{h,k},a^{\prime}_{h,k})_{h=0}^{\infty}$ we consider the design points $x_{k}:=\phi(\tau_{k})-\phi(\tau^{\prime}_{k})=\sum_{h=0}^{\infty}\gamma^{h}(\phi(s_{h,k},a_{h,k})-\phi(s^{\prime}_{h,k},a^{\prime}_{h,k}))$ and the preference labels $y_{k}=\mathbbm{1}(\tau_{k}\succ\tau^{\prime}_{k})$. In round $t$, the preference dataset is $\mathcal{D}_{t}=\{(x_{k},y_{k})\}_{k=1}^{t-1}$ and the corresponding (constrained) maximum likelihood estimator (MLE) is given by $\hat{\theta}_{t}=\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq B}\mathcal{L}_{\mathcal{D}_{t}}(\theta)$, where
@@
is the negative log-likelihood of the Bradley-Terry model (1). The loss function (2) is the familiar logistic loss from logistic regression (shalev2014understanding). In particular, it is a convex problem that can be solved efficiently using standard methods such as LBFGS (liu1989limited). Moreover, we have the following time-uniform confidence result. {restatable}lemmaconfidence Let $\lambda\geq 0$ and define the design matrix at time $t$ given by $V_{t}=\lambda I+\sum_{k=1}^{t-1}x_{k}x_{k}^{\top}$. Then, with probability $1-\delta$, for all $t\in\mathbb{N}$, the true reward parameter $\theta^{*}$ is contained in the ellipsoid
$\mathcal{E}_{t}(\delta):=\left\{{\theta:\left\lVert\theta-\hat{\theta}_{t}\right\rVert_{V_{t}}^{2}\leq\beta_{t}(\delta)^{2}:=\mathcal{O}\left({\kappa\left[{\log\left({\frac{1}{\delta}}\right)+d\log\left({\frac{t-1}{d}}\right)}\right]+\lambda}\right)}\right\}.$
Here, $\kappa:=\max_{\theta\in\mathcal{B}^{d}(B),x\in\mathcal{B}^{d}(2LH_{\gamma})}1/\dot{\sigma}\left({\langle\theta,x\rangle}\right)$ denotes the Lipschitz constant of the inverse sigmoid function, and $H_{\gamma}=(1-\gamma)^{-1}$ the effective horizon of the MDP. The above lemma hinges on a result for likelihood-ratio confidence sets by lee2024unified. The proof and the precise constants are deferred to Appendix D.
@@
Compared to the standard analysis of stochastic linear bandits abbasi2011improved, our parameter $\beta_{t}$ includes an additional factor of $\sqrt{\kappa}$, which arises naturally due to preference-based feedback. This result improves upon the bound provided by zhu2023principled, which incurs a larger factor of $\kappa$ instead of $\sqrt{\kappa}$. While the $\sqrt{\kappa}$ factor can theoretically be avoided by constructing confidence sets using the Hessian of the negative log-likelihood $\mathcal{L}_{\mathcal{D}_{t}}$ (lee2024unified; das2024active), it reappears in the regret bounds as shown in (das2024active). We adopt confidence sets based on $V_{t}$, as this facilitates preference-free exploration, and both $V_{t}$ and its inverse can be efficiently updated via rank-one operations, unlike Hessian-based approaches.
@@
Many approaches to regret minimization and pure exploration in RLHF rely on maximizing an exploration bonus of the form $\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}$ or $\mathbb{E}_{\tau\sim\mathbb{P}_{\pi_{t}},\tau^{\prime}\sim\mathbb{P}_{\pi_{t}^{\prime}}}\left\lVert\phi(\tau)-\phi(\tau^{\prime})\right\rVert_{V_{t}^{-1}}^{2}$. Although such methods yield provable guarantees for regret (pacchiano2021dueling) or last-iterate suboptimality (das2024active), they are computationally intractable in RL settings (see Appendix E.2). To address this, we adopt a randomized exploration scheme inspired by Thompson sampling algorithms for linear bandits. In line with pmlr-v54-abeille17a, we sample the reward parameter from an inflated version of the confidence set defined in Lemma 2, which produces a computationally efficient alternative to optimism-based approaches. Furthermore, we also show that this randomized strategy extends to the pure exploration setting.
@@
We assume access to an $(\varepsilon,\delta)$-PAC oracle, $\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}$, for the RL problem. That is, a polynomial-time algorithm that produces for every $\varepsilon>0$, $\delta\in(0,1)$, and $\theta\in\mathbb{R}^{d}$ a policy $\pi=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\theta,\varepsilon,\delta)$ such that with probability at least $1-\delta$ we have $V^{*}_{\theta}-V_{\theta}^{\pi}\leq\varepsilon$.
This assumption of a PAC-RL oracle is satisfied in several settings, including tabular and linear MDPs (dann2019policy; menard2021fast; al2021navigating; wagenmaker2022beyond; he2021uniform). Moreover, it is relatively mild compared to stronger oracles considered in the RLHF literature (zhan2024provablereward), such as reward-free algorithms (wang2020reward; kaufmann2021adaptive; menard2021fast). In practice, common choices for $\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}$ are policy optimization methods such as proximal policy optimization (PPO) (schulman2017proximal) or soft actor critic (haarnoja2018soft), which have shown strong empirical performance in continuous control and large-scale applications such as training large language models.
@@
Our algorithm randomized preference optimization (RPO) presented below comes in two variants: (i) RPO-Regret (Algorithm 1) which balances exploration and exploitation for regret minimization, and (ii) RPO-Explore (Algorithm 2) which performs preference-free exploration and collects a single batch of preferences at the end. In RPO-Regret, we sample in each round a reward parameter from a confidence set (see Lemma 2) inflated by $\sqrt{d}$, compute the policy $\pi_{t}$ via the RL oracle, and update the reward estimate by maximum likelihood using the newly collected preference. In contrast, RPO-Explore samples reward parameters centered at zero, stores trajectory pairs during exploration, and collects a single batch of preferences at the end.
Input: $T\in\mathbb{N},\ \delta\in(0,1),\ \lambda>0$
1 Initialize: $\varepsilon=1/\sqrt{T}$; $\delta^{\prime}=\delta/5$; $V_{1}=\lambda I$; $\mathcal{D}_{1}=\emptyset$; $\hat{\theta}_{1}=0$; and choose $\pi_{0}$.
2for $t=1,2,\dots,T$ do
 $\tilde{\theta}_{t}\sim\mathcal{N}\big(\hat{\theta}_{t},\beta_{t}(\delta^{\prime})^{2}V_{t}^{-1}\big)$